Show algebraically that segment A (2,0) A' (0,2) is parallel to segment C (8,12) C' (12,8)
I dont understand that question. Could someone help with this? Thank you.

Answer by stanbon(75887)

Show algebraically that segment A (2,0) A' (0,2) is parallel to segment C (8,12) C' (12,8)
-------------------------
Check the slopes:
AA': (2-0)/(0-2) = 2/-2 = -1
---------------------------------------
CC': (8-12)/(12-8) = -4/4 = -1
------------------------------------
Since the slopes are the same, the lines are parallel.
Cheers,
Stan H.